### v2

@@ 1 Introduction
Figure 1: (0(a)) A symmetric MBB beam design domain. (0(b)) A schematic for node (blue texts) and element (red texts) numbering scheme [1]. (0(c)) Multiple load cases domain. Texts in circles indicate load cases. (0(d)) Domain with passive regions
@@ 3.1 Element Connectivity and Nodal Coordinates matrices (lines 15-28)
Lines 3-8 of PyHexTop are needed to install all the required libraries. Let HNex and HNey be the number of hexagonal elements in $x$ and $y$ directions, respectively. Each element consists of six nodes. Each node contains two degrees of freedom (DOFs). The Numpy array NstartVs (line 15) assigns node numbers of elements row by row, starting from the bottom (See 0(b)). DOFstartVs (line 16) stores the DOFs for the symmetrical half quadrilaterals of each hexagonal element, and the order of storage follows the numbering of the nodes, with each node’s DOFs arranged sequentially. This array, modified for DOF organization, is horizontally replicated eight times and then vertically replicated $2\times\texttt{HNey}\times\texttt{HNex}$ times. The resulting matrix is named NodeDOFs (line 17). The redundant DOFs related to symmetry are removed using the setdiff Numpy function, and the resulting matrix is stored in ActualDOFs (line 18). The final connectivity matrix HoneyDOFs (line 19) contains organized DOFs relevant to the honeycomb structure, with the specified columns from ActualDOFs. Readers are recommended to refer to [1] for more detail.
@@ 3.4 MBB beam optimized results
Figure 3: Optimized results: multi-load cases (2(a)) 2-load cases (2(b)) 4-load cases; Passive regions (2(c)) $\mathtt{ft=1}$ (2(d)) $\mathtt{ft=2}$
@@ 4.1 Multiple loads
The optimized designs for 2-load and 3-load cases are depicted in Figs. 2(a) and 2(b).
@@ 4.2 Passive design domains
The code is called as below and optimized designs for ft=1 and ft=2 are displayed in Figs. 2(c) and 2(d), respectively.

### v3

@@ 1 Introduction
Figure 1: (1(a)) A symmetric MBB beam design domain. (1(b)) A schematic for node (blue texts) and element (red texts) numbering scheme [1]. (1(c)) Multiple load cases domain. Texts in circles indicate load cases. (1(d)) Domain with passive regions
@@ 3.1 Element Connectivity and Nodal Coordinates matrices (lines 15-28)
Lines 3-8 of PyHexTop are needed to install all the required libraries. Let HNex and HNey be the number of hexagonal elements in $x$ and $y$ directions, respectively. Each element consists of six nodes. Each node contains two degrees of freedom (DOFs). The Numpy array NstartVs (line 15) assigns node numbers of elements row by row, starting from the bottom (See 1(b)). DOFstartVs (line 16) stores the DOFs for the symmetrical half quadrilaterals of each hexagonal element, and the order of storage follows the numbering of the nodes, with each node’s DOFs arranged sequentially. This array, modified for DOF organization, is horizontally replicated eight times and then vertically replicated $2\times\texttt{HNey}\times\texttt{HNex}$ times. The resulting matrix is named NodeDOFs (line 17). The redundant DOFs related to symmetry are removed using the setdiff Numpy function, and the resulting matrix is stored in ActualDOFs (line 18). The final connectivity matrix HoneyDOFs (line 19) contains organized DOFs relevant to the honeycomb structure, with the specified columns from ActualDOFs. Readers are recommended to refer to [1] for more detail.
@@ 3.4 MBB beam optimized results
Figure 3: Optimized results: multi-load cases (3(a)) 2-load cases (3(b)) 4-load cases; Passive regions (3(c)) $\mathtt{ft=1}$ (3(d)) $\mathtt{ft=2}$
@@ 4.1 Multiple loads
The optimized designs for 2-load and 3-load cases are depicted in Figs. 3(a) and 3(b).
@@ 4.2 Passive design domains
The code is called as below and optimized designs for ft=1 and ft=2 are displayed in Figs. 3(c) and 3(d), respectively.